25.0 g H, are reacted, how much Fe is

produced?

Molar masses: 2.02 %mol Ha, 55.9 %/mol Fe

A. 0.903 g Fe

B. 1,112 g Fe

C. 692 g Fe

D. 346 g Fe


1
Expert's answer
2022-05-31T16:16:03-0400

Fe2O3 + 3H2 = 2Fe + 3H2O


m(Fe) = (m(H2)•2Mr(Fe)) / 3Mr(H2)

m(Fe) = (25 g • 2•55.9 g/mol) / 3•2.02 g/mol = 461.2 g


m(Fe) = 25•55.9/2.02 = 692 g